After 2238.6 years, how much of an original
1.15 g of radium-226 remains? Its half-life is
1599 years.
1
Expert's answer
2019-09-18T07:00:51-0400

We need to know how mamy half-life periods arę passed

n = 2238.6/1599 = 1.4

Then we have the equation:

0.51.4= x

Where x is percent of radium-226 remains

x = 0.379

Then

m = m0*x = 1.15*0.379 = 0.436 g